MG770- Financial Reporting, Financial Statement Analysis, and Valuation
Homework Assignment
Week # 4
__________________________________________________________________________

Using the following ratios to analyze a company financial performance.

Profitability Analysis
Profit Margin Analysis
Asset Turnover
Fixed Asset Turnover
Rate of Return on Common Shareholders’ Equity
Earning Per Common Share

Instruction:

1] Please search a company (example as Google, facebook, Macy’s and others) with a set of current financial statements and apply the ratios listed above to analyze a company financial performance

2] Please provide the interpretation on each ratio analysis reflecting a company current financial situation.
Answer #1

Solution:

Financial ratios are very important measures to analyse the performance of a company. There are four major categories of financial ratios-Profitability, Efficiency, Liquidity and solvency ratio.

The chosen company is Google and year ending 2017 data is taken from Yahoo Finance.

Question 1

1. Profitability analysis

Profit Margin : Profit Margin = Net profit / Sales

Net Profit = 12.662 Billion , Sales = 110.855 Billion

Profit Margin = 12.662/110.885 = 11.42%

2. Total asset turnover = Sales / Total Asset

Total Asset = 197.295 Billion

Total Asset Turnover = 110.855 / 197.295 = 0.56

3. Fixed Asset turnover = Sales / Fixed Assets

Fixed Assets = 72.987

Fixed Asset turnover = 110.855/ 72.987 = 1.52

4. Rate of return on shareholder's equity or ROE

ROE = Net income / Equity

Equity = 152.502 Billion

ROE = 12.662/152.502 = 8.30%

5. Earning per share = Earning (Net Income) / Number of share

Number of share = 349.615 Million

Earning per share = 12,662 Million / 349.615 = 36.21

Question 2.

Interpretation of each ratio

1. Profit Margin : Profit margin is 11.42%, and it is very good considering the fact that Google has now become a matured company.

2. Total asset turnover ratio is 0.56 and it means that Google is able to generate 0.56 dollar of sales for every dollar of asset. This is again a good ratio though this can be improved considering the fact that Google is not in the asset heavy industry

3. Fixed asset turnover is 1.52 and it means that Google generates 1.52 Dollar of revenue for every dollar of fixed assets. This is good as it implies that the company has more current assets than fixed assets.

4. ROE is 8.3% and it implies that the company generates 8.3% profit for given total equity. This number is slightly lower due to the fact that the company has high retained earnings

5. EPS is 36.21, EPS means the company is able to generates 36.21 Dollars of profit per share. This is slightly less comparing the share price of Google that is around $1077. It makes the PE ratio as 1077/36.21 = 29.73